# M-learner:A Flexible And Powerful Framework To Study Heterogeneous Treatment Effect In Mediation Model

### v1

@@ Abstract
We propose a novel method, termed the M-learner, for estimating heterogeneous indirect and total treatment effects and identifying relevant subgroups within a mediation framework. The procedure comprises four key steps. First, we compute individual-level conditional average indirect/total treatment effect Second, we construct a distance matrix based on pairwise differences. Third, we apply t-SNE to project this matrix into a low-dimensional Euclidean space, followed by K-means clustering to identify subgroup structures. Finally, we calibrate and refine the clusters using a threshold-based procedure to determine the optimal configuration. To the best of our knowledge, this is the first approach specifically designed to capture treatment effect heterogeneity in the presence of mediation. Experimental results validate the robustness and effectiveness of the proposed framework. Application to the real-world Jobs II dataset highlights the broad adaptability and potential applicability of our method.Code is available at https://anonymous.4open.science/r/M-learner-C4BB.
@@ 1 Introduction
For heterogeneous total treatment effects, Foster et al. [2011] introduced the Virtual Twins to estimate heterogeneous total treatment effects with RCT data. Several methodological approaches within observational studies enable the estimation of flexible and accurate models of heterogeneous total treatment effects. Shalit et al. [2017], Shi et al. [2019], Johansson et al. [2016], Hassanpour and Greiner [2019] use neural networks to estimate heterogeneous total treatment effects. Athey and Imbens [2016], Wager and Athey [2018], Athey and Wager [2019] use random forests to estimate heterogeneous total treatment effects. Künzel et al. [2019] proposed the Meta-learners which consist of S-learner, T-learner, and X-learner. Nie and Wager [2021] proposed R-learner, Zhang et al. [2022] extented the R-learner to continuous treatment scenario. And other methods have been proposed, such as DR-learner and Lp-R-learner [Kennedy, 2023]. Dwivedi et al. [2020] proposed model selection solution in causal inference.
When a mediator is present, the primary analytical focus is on evaluating the indirect treatment effect on the outcome via the mediator. These method decompose the average treatment effect into direct treatment effect and indirect effect[Lin et al., 1997, Preacher, 2015, Robins and Greenland, 1992, Petersen et al., 2006, van der Laan and Petersen, 2008, Imai et al., 2010, Tchetgen and Shpitser, 2012, VanderWeele, 2015, Vandenberghe et al., 2017, Dorresteijn et al., 2011, Pearl, 2022, VanderWeele and Vansteelandt, 2009, Vansteelandt and VanderWeele, 2012, Angrist, 2004, Imbens, 2004]. Recently, to address commonly observed intermediated confounders that would be affected by the covariates and then affect both mediators and outcome, multiple methods have been developed to extend the classical metiation analysis[Tchetgen and VanderWeele, 2014, Díaz et al., 2021, 2023, Gilbert et al., 2024]. Ge et al. [2023, 2025], Luo et al. [2025] and Wang and Song [2025]proposed the method to use reinforcement learning to deal with the dynamic mediation analysis. Cheng et al. [2022] use deep learning to estimate causal effects in mediation model. However, these methods do not account for treatment effect heterogeneity in mediation model. In contrast, our approach is the first to incorporate heterogeneity of treatment effects within a mediation framework. This enables the identification of subgroups characterized by distinct indirect treatment effects and provides a principled way to evaluate the effectiveness of the mediator and causal decision making.
@@
Our approach facilitates the identification of subgroups exhibiting differential response patterns through mediated pathways, thereby improving the interpretability and clinical relevance of treatment effect estimation.
The key contributions of this paper are:
@@ 2.1 Model Formulation
where $X_{i}\in\mathbb{R}^{d}$ is a d-dimensional covariate or feature vector, $W_{i}\in\{0,1\}$ is the treatment-assignment indicator,$Y_{i}(0,M_{i}(0))\in\mathbb{R}$ is the potential outcome of unit $i$ when $i$ is assigned to control group, and the mediator $M$ is assigned to control group, $Y_{i}(0,M_{i}(1))\in\mathbb{R}$ is the potential outcome of unit $i$ when $i$ is assigned to control group, and the mediator $M$ is assigned to treatment group, $Y_{i}(1,M_{i}(0))\in\mathbb{R}$ is the potential outcome of unit $i$ when $i$ is assigned to treatment group, while the mediator $M$ is assigned to control group,$Y_{i}(1,M_{i}(1))\in\mathbb{R}$ is the potential outcome of unit $i$ when $i$ is assigned to treatment group, and the mediator $M$ is assigned to treatment group. What’s more, this paper follows the four assumptions specified in the VanderWeele and Vansteelandt [2009] regarding the interrelations among covariate $X$, outcome $Y$, and the mediator $M$. When a mediator $M$ is present, the relationships among $X$, $Y$, and $M$ follow the structure shown in the Directed Acyclic Graph (DAG) in Figure 1 (b). The treatment effect can be transmitted through two distinct pathways. If the treatment effect is transmitted along path $a\rightarrow b$ via the mediator $M$, this is referred to as the indirect treatment effect (ITE). If the treatment effect is transmitted directly to $Y$ along path $c$, we refer to it as the direct treatment effect (DTE). The total treatment effect (TTE) is the combined influence of both the direct and indirect pathways. These effects are formally defined as follows,
$ITE$ $=$ $Y(1,M(1))-Y(0,M(1)),DTE=Y(0,M(1))-Y(0,M(0)),$
@@
Figure 1: (a)Directed Acyclic Graph between the covariates $X$ and response $Y$, (b)Directed Acyclic Graph between the mediator $M$, covariates $X$ and response $Y$,(c)Pipeline of the M-learner.
@@
For a new unit $i$ with covariate vector $x_{i}$, to decide whether to give the treatment via total treatment effect, we wish to estimate the individual total treatment effect (ITTE) of each $i,D_{i}$, which is defined as $D_{i}:=Y_{i}(1,M(1))-Y_{i}(0,M(0)).$ However, we can not observe $D_{i}$ for any unit, and $D_{i}$ is not identifiable without strong assumptions. Instead, Künzel et al. [2019] proposed that we can estimate the CATTE function, which is defined as $\tau^{TOT}(x)=\mathbb{E}[D|X=x]=\mathbb{E}[Y(1,M(1)-Y(0,M(0)))|X=x],$ the best estimator for the CATEE is also the best estimator for the ITTE in terms of the mean squared error (MSE). Let $\hat{\tau}_{i}^{TOT}$ be an estimator of $D_{i}$ and decompose the MSE at $x_{i}$,
$\mathbb{E}[(D_{i}-\hat{\tau}_{i}^{TOT})^{2}|X_{i}=x_{i}]$ $=$ $\mathbb{E}[(D_{i}-\tau^{TOT}(x_{i}))^{2}|X_{i}=x_{i}]+\mathbb{E}[(\tau^{TOT}(x_{i})-\hat{\tau}_{i}^{TOT})^{2}].$ (1)
Because we can not affect the first term in 1, the estimator that minimizes the MSE for the ITTE of $i$ also minimizes the MSE for the CATTE at $x_{i}$. To decide whether to give the treatment via indirect treatment effect, we wish to estimate the individual indirect treatment effect (IITE) of each $i$, where the definition of IITE is $Y_{i}(1,M(1))-Y_{i}(1,M(0))|X]$. Analogous to Expression 1, the CAITE can also be regarded as the optimal estimator of the IITE under the MSE criterion, where the definition of CAITE is $\tau^{ITE}:=\mathbb{E}[Y(1,M(1))-Y(1,M(0))|X].$ With the estimation of CATTE and CAITE, we can use the estimation to study the subgroups of TTE and ITE. For subgroups via TTE, it can be defined as $\mathbb{E}[Y(1,M(1))-Y(0,M(0))\in\mathcal{U}_{i}^{TTE}|X],i=1,2,3,\cdots,$ where $\mathcal{U}_{i}^{TTE}$ is the set, the number of the sets are unknown, and when there is no heterogeneity, the number of the set is 1. For subgroups via ITE, it can be defined as $\mathbb{E}[Y(1,M(1))-Y(1,M(0))\in\mathcal{U}_{i}^{ITE}|X],i=1,2,3,\cdots,$ where $\mathcal{U}_{i}^{ITE}$ is the set, the number of the sets are unknown, and when there is no heterogeneity, the number of the set is 1.
@@ 2.2 M-learner
We propose an algorithm called M-learner [^1] We refer to the method as the M-learner, as it is specifically designed to operate within the framework of mediation models.shown in Figure 1 (c) to address the aforementioned problem. This algorithm consists of three main steps:
@@
[…]
@@ 3.1 Subgroup analysis without mediators
These results indicate that when the underlying heterogeneity is relatively simple, both RF and XGB perform well in identifying heterogeneous treatment groups. As the heterogeneity structure increases in complexity, the performance of the proposed method deteriorates. However, as the complexity of the heterogeneity structure increases, XGB consistently outperforms RF. Moreover, our proposed method not only effectively identifies heterogeneous treatment groups but also maintains a low Type I error rate, underscoring its robustness and reliability across a variety of conditions. We compared our algorithm with K-Means, and the results show that K-Means lacks the ability to identify heterogeneous groups ( Appendix A.4 provides detailed procedures and results of the comparative algorithms can be found in Appendix A.4 Table A.12,A.13 and Figure A.9).
@@ 3.2 Subgroup analysis with a mediator
To identify subgroups that benefit from the treatment effects through the mediator, we design seven scenarios that reflect varying real-world heterogeneity structures . Specifically, the scenarios represent (i) simple heterogeneity, all treatment effect via mediator (Simple-All); (ii) simple heterogeneity,part treatment effect via mediator (Simple-Part); (iii) complex heterogeneity,all treatment effect via mediator (Complex-Part); (iv) complex heterogeneity, part treatment effect via mediator (Complex-Part); (v) simple no heterogeneity, no treatment effect via mediator (Simple-Null1); (vi) simple no heterogeneity, no treatment effect via mediator (Simple-Null2); (vii) simple no heterogeneity, all units benefit from the treatment and all treatment effect via mediator (Simple-Global); further experimental details are available in the Appendix A.5.2.
@@
In scenarios with a simple underlying structure, both learners are capable of effectively identifying subgroups reflecting treatment effects transmitted through the mediator. However, under more complex settings, XGBoost consistently outperforms Random Forest. Specifically, the M-learner framework with XGBoost demonstrates strong capacity to detect indirect treatment effects and to recover meaningful subgroups accordingly. It also robustly rejects cases where the indirect effect is zero, indicating an ineffective mediator. Simple-Null1 and Simple-Null2 both assume a null indirect effect, but differ in whether the mediator influences the outcome. In Simple-Null1, the treatment does not affect the mediator, while in Simple-Null2, the mediator has no association with the outcome, implying it is not a true mediator. Our proposed method successfully identifies both types of scenarios using either base learner. For non-heterogeneous settings, we simulate only simple scenarios, while in heterogeneous cases, we consider both simple and complex structures to reflect realistic uncertainties regarding the complexity of heterogeneity. Simple scenarios serve as a baseline to filter out non-heterogeneous structures. Overall, the experimental results strongly support our hypotheses: across varying degrees of complexity, the proposed approach reliably determines the effectiveness of the mediator $M$, estimates the mediation effect, and identifies the corresponding subgroups. Figure 2 (c) and (d) show the ECDFs of p-values under different scenarios. The figure elucidates why the $p_{leaf}$ is instrumental in discerning non-heterogeneous scenarios and in assessing the efficacy of a mediator. We compared our algorithm with K-Means, and the results show that K-Means lacks the ability to identify heterogeneous groups in the mediation model(the Appendix A.4 provides detailed procedures and results of the comparative algorithms can be found in Appendix Table A.14,A.15 ,A.16 and Figure A.9).
3.3 Effect of sample size on M-learner performance and visualization of the M-learner
@@
To assess the performance of the M-learner under varying sample sizes, we additionally compare the results to those obtained with a sample size of $500$, while holding all other conditions constant. In Appendix A.3, the Table A.7-A.11 and the Figure A.4-A.8 show the detailed results. The results indicate that, in the absence of mediators, the M-learner is capable of accurately identifying heterogeneous treatment regions in the simple scenario, even with a reduced sample size. However, its performance deteriorates substantially in the complex scenario, suggesting that the identification of treatment effect heterogeneity under more complex conditions requires larger sample sizes. Notably, the M-learner continues to perform well in detecting non-heterogeneous scenarios, even when the available sample size is limited.
@@
In the presence of a mediator, the X-learner remains effective in identifying heterogeneous treatment regions mediated by the mediator, even with a reduced sample size. In the complex scenario, the performance of the RF-based learner declines significantly, whereas the performance degradation of the XGB-based learner is relatively limited. This suggests that, under limited sample sizes, XGB serves as a more robust base learner. The M-learner also demonstrates stable performance; however, accurately estimating complex relationships among $M$, $X$, and $Y$ still requires a sufficiently large sample size to reliably identify heterogeneous treatment effects.
@@
Comparisons of ECDFs across varying sample sizes and scenarios reveal that, in the simple scenario, the ECDFs for sample sizes of $1000$ and $500$ exhibit minimal divergence. In contrast, more pronounced differences are observed in the complex scenario. Notably, an ECDF curve positioned closer to the upper-left corner reflects stronger treatment effect heterogeneity, whereas a curve nearer to the lower-right corner indicates weaker heterogeneity. A comprehensive comparison across different learners, scenarios, and sample sizes is presented in Appendix Figure A.5-A.8.
@@
The visualization results of the M-learner are provided in Appendix A.5 for reference.

### v3

@@ Abstract
We propose a novel method, termed the M-learner, for estimating heterogeneous indirect and total treatment effects and identifying relevant subgroups within a mediation framework. The procedure comprises four key steps. First, we compute individual-level conditional average indirect/total treatment effect Second, we construct a distance matrix based on pairwise differences. Third, we apply t-SNE to project this matrix into a low-dimensional Euclidean space, followed by K-means clustering to identify subgroup structures. Finally, we calibrate and refine the clusters using a threshold-based procedure to determine the optimal configuration. To the best of our knowledge, This is the first method capable of revealing the complex relationships among X, M, and Y within mediation analysis while effectively controlling the Type I error rate. Experimental results validate the robustness and effectiveness of the proposed framework. Application to the real-world Jobs II dataset highlights the broad adaptability and potential applicability of our method. Code is available at https://anonymous.4open.science/r/M-learner-C4BB.
@@ 1 Introduction
For heterogeneous total treatment effects, Foster et al. [2011] introduced the Virtual Twins to estimate heterogeneous total treatment effects with RCT data. Several methodological approaches within observational studies enable the estimation of flexible and accurate models of heterogeneous total treatment effects. Shalit et al. [2017], Shi et al. [2019], Johansson et al. [2016], Hassanpour and Greiner [2019] use neural networks to estimate heterogeneous total treatment effects. Athey and Imbens [2016], Wager and Athey [2018], Athey and Wager [2019] use random forests to estimate heterogeneous total treatment effects. Künzel et al. [2019] proposed the Meta-learners which consist of S-learner, T-learner, and X-learner. Nie and Wager [2021] proposed R-learner, Zhang et al. [2022] extented the R-learner to continuous treatment scenario. And other methods have been proposed, such as DR-learner and Lp-R-learner [Kennedy, 2023]. Dwivedi et al. [2020] proposed model selection solution in causal inference. Kim et al. [2024] proposed the causal clustering method.
When a mediator is present, the primary analytical focus is on evaluating the indirect treatment effect on the outcome via the mediator. These method decompose the average treatment effect into direct treatment effect and indirect effect[Lin et al., 1997, Preacher, 2015, Robins and Greenland, 1992, Petersen et al., 2006, van der Laan and Petersen, 2008, Imai et al., 2010, Tchetgen and Shpitser, 2012, VanderWeele, 2015, Vandenberghe et al., 2017, Dorresteijn et al., 2011, Pearl, 2022, VanderWeele and Vansteelandt, 2009, Vansteelandt and VanderWeele, 2012, Angrist, 2004, Imbens, 2004]. Recently, to address commonly observed intermediated confounders that would be affected by the covariates and then affect both mediators and outcome, multiple methods have been developed to extend the classical metiation analysis[Tchetgen and VanderWeele, 2014, D´ıaz et al., 2021, 2023, Gilbert et al., 2024]. Ge et al. [2023, 2025], Luo et al. [2025] and Wang and Song [2025]proposed the method to use reinforcement learning to deal with the dynamic mediation analysis. Cheng et al. [2022] use deep learning to estimate causal effects in mediation model. However, these methods do not account for treatment effect heterogeneity in mediation model. Recently, Ting and Linero [2025] proposed the method to use BART to estimate the heterogeneous mediation effect. This method is primarily designed for estimating heterogeneous treatment effects and does not provide an effective solution for determining the existence of heterogeneity or for identifying heterogeneous subgroups. In contrast, our method is the first to incorporate treatment effect heterogeneity within a mediation framework and to provide an approach for detecting the existence of heterogeneous subgroups as well as heterogeneous regions. It can effectively identify heterogeneity and the corresponding region. This enables the identification of subgroups characterized by distinct indirect treatment effects and provides a principled way to evaluate the effectiveness of the mediator and causal decision making.
@@
Our approach facilitates the identification of subgroups exhibiting differential response patterns through mediated pathways, thereby improving the interpretability and clinical relevance of treatment effect estimation. The key contributions of this paper are:
@@ 2.1 Model Formulation
[…]
$ITE$ $=$ $Y(1,M(1))-Y(1,M(0)),DTE=Y(1,M(0))-Y(0,M(0)),$
@@
Figure 1: (a)Directed Acyclic Graph between the covariates $X$ and response $Y$, (b)Directed Acyclic Graph between the mediator $M$, covariates $X$ and response YY,(c)Pipeline of the M-learner.
@@
For a new unit $i$ with covariate vector $x_{i}$, to decide whether to give the treatment via total treatment effect, we wish to estimate the individual total treatment effect (ITTE) of each $i,D_{i}$, which is defined as $D_{i}:=Y_{i}(1,M(1))-Y_{i}(0,M(0)).$ However, we can not observe $D_{i}$ for any unit, and $D_{i}$ is not identifiable without strong assumptions. Instead, Künzel et al. [2019] proposed that we can estimate the CATTE function, which is defined as $\tau^{TOT}(x)=\mathbb{E}[D|X=x]=\mathbb{E}[Y(1,M(1)-Y(0,M(0)))|X=x],$ the best estimator for the CATTE is also the best estimator for the ITTE in terms of the mean squared error (MSE). Let $\hat{\tau}_{i}^{TOT}$ be an estimator of $D_{i}$ and decompose the MSE at $x_{i}$,
$\mathbb{E}[(D_{i}-\hat{\tau}_{i}^{TOT})^{2}|X=x_{i}]$ $=$ $\mathbb{E}[(D_{i}-\tau^{TOT}(x_{i}))^{2}|X=x_{i}]+\mathbb{E}[(\tau^{TOT}(x_{i})-\hat{\tau}_{i}^{TOT})^{2}].$ (1)
Because we can not affect the first term in 1, the estimator that minimizes the MSE for the ITTE of $i$ also minimizes the MSE for the CATTE at $x_{i}$. To decide whether to give the treatment via indirect treatment effect, we wish to estimate the individual indirect treatment effect (IITE) of each $i$, where the definition of IITE is $Y_{i}(1,M(1))-Y_{i}(1,M(0))$. Analogous to Expression 1, the CAITE can also be regarded as the optimal estimator of the IITE under the MSE criterion, where the definition of CAITE is $\tau^{ITE}:=\mathbb{E}[Y(1,M(1))-Y(1,M(0))|X].$ With the estimation of CATTE and CAITE, we can use the estimation to study the subgroups of TTE and ITE. For subgroups via TTE, it can be defined as $\mathbb{E}[Y(1,M(1))-Y(0,M(0))\in\mathcal{U}_{i}^{TTE}|X],i=1,2,3,\cdots,$ where $\mathcal{U}_{i}^{TTE}$ is the set, the number of the sets are unknown, and when there is no heterogeneity, the number of the set is 1. For subgroups via ITE, it can be defined as $\mathbb{E}[Y(1,M(1))-Y(1,M(0))\in\mathcal{U}_{i}^{ITE}|X],i=1,2,3,\cdots,$ where $\mathcal{U}_{i}^{ITE}$ is the set, the number of the sets are unknown, and when there is no heterogeneity, the number of the set is 1.
@@ 2.2 M-learner
We propose an algorithm called M-learner [^1] We refer to the method as the M-learner, as it is specifically designed to operate within the framework of mediation models.shown in Figure 1 (c) to address the aforementioned problem. This algorithm consists of four main steps:
@@
First, we evaluate the treatment effect difference between each pair of units $i$ and $j$, which we refer to as the distance. It is defined by $dis(i,j)$, The distance metric can be defined as Euclidean distance, Manhattan distance, or other formulations. In the following analysis, we consistently adopt Euclidean distance to compute the distance between units $i$ and $j$ which means $dis(i,j)=(\hat{\tau}_{i}-\hat{\tau}_{j})^{2}$. Consequently, an $n\times n$ distance matrix is obtained. This matrix is then projected into a two-dimensional Euclidean space using t-SNE[Van der Maaten and Hinton, 2008], where each unit $i$ is assigned a coordinate. Subsequently, K-means clustering is performed on the projected points in the Euclidean space. The range of cluster numbers (from 2 to $k$) for K-means clustering must be specified in advance, based on the sample size and the number of covariates; an excessively large $k$ can result in substantial overfitting, based on our experience, we recommend $k=\left\lfloor\sqrt{d}\right\rfloor+2$, where $\lfloor\rfloor$ represents rounding down to the nearest integer. Then, a decision tree is employed to model the clustering results obtained for each predefined number of clusters, with the objective of mapping the unknown categories to interpretable categorical information. For each leaf of the decision tree, we refer to it as a subtype. We use $p_{leaf}$to select the final, unique subtype grouping from different decision tree results. The definition of $p_{leaf}$is as follows,
@@ 3.1 Subgroup analysis without mediators
Table 1: The table summarizes the correct covariates in profiles in each scenario. $X_{1}/X_{2}$:Final profile contain $X_{1}/X_{2}$, $X_{1}\&X_{2}$: final profile contain both $X_{1}$ and $X_{2}$$X_{1}$ and $X_{2}$ are the two variables associated with treatment effect heterogeneity. Each value denotes the count of occurrences across 100 simulations.
@@
These results indicate that when the underlying heterogeneity is relatively simple, both RF and XGB perform well in identifying heterogeneous treatment groups. As the heterogeneity structure increases in complexity, the performance of the proposed method deteriorates. However, as the complexity of the heterogeneity structure increases, XGB consistently outperforms RF. Moreover, our proposed method not only effectively identifies heterogeneous treatment groups but also maintains a low Type I error rate, underscoring its robustness and reliability across a variety of conditions. We compared our algorithm with K-Means, and the results show that K-Means lacks the ability to identify heterogeneous groups (Appendix A.5 provides detailed procedures and results of the comparative algorithms can be found in Appendix A.5 Table A.12,A.13 and Figure A.10, Figure A.4 shows the boundary of the selected heterogeneous treatment region).
@@ 3.2 Subgroup analysis with a mediator
To identify subgroups that benefit from the treatment effects through the mediator, we design seven scenarios that reflect varying real-world heterogeneity structures . Specifically, the scenarios represent (i) simple heterogeneity, all treatment effect via mediator (Simple-All); (ii) simple heterogeneity,part treatment effect via mediator (Simple-Part); (iii) complex heterogeneity,all treatment effect via mediator (Complex-All); (iv) complex heterogeneity, part treatment effect via mediator (Complex-Part); (v) simple no heterogeneity, no treatment effect via mediator (Simple-Null1); (vi) simple no heterogeneity, no treatment effect via mediator (Simple-Null2); (vii) simple no heterogeneity, all units benefit from the treatment and all treatment effect via mediator (Simple-Global); further experimental details are available in the Appendix A.10.2.
@@
In scenarios with a simple underlying structure, both learners are capable of effectively identifying subgroups reflecting treatment effects transmitted through the mediator. However, under more complex settings, XGBoost consistently outperforms Random Forest. Specifically, the M-learner framework with XGBoost demonstrates strong capacity to detect indirect treatment effects and to recover meaningful subgroups accordingly. It also robustly rejects cases where the indirect effect is zero, indicating an ineffective mediator. Simple-Null1 and Simple-Null2 both assume a null indirect effect, but differ in whether the mediator influences the outcome. In Simple-Null1, the treatment does not affect the mediator, while in Simple-Null2, the mediator has no association with the outcome, implying it is not a true mediator. Our proposed method successfully identifies both types of scenarios using either base learner. For non-heterogeneous settings, we simulate only simple scenarios, while in heterogeneous cases, we consider both simple and complex structures to reflect realistic uncertainties regarding the complexity of heterogeneity. Simple scenarios serve as a baseline to filter out non-heterogeneous structures. Overall, the experimental results strongly support our hypotheses: across varying degrees of complexity, the proposed approach reliably determines the effectiveness of the mediator $M$, estimates the mediation effect, and identifies the corresponding subgroups. Figure 2 (c) and (d) show the ECDFs of p-values under different scenarios. The figure elucidates why the $p_{leaf}$ is instrumental in discerning non-heterogeneous scenarios and in assessing the efficacy of a mediator. Appendix Figure A.4 shows the selected heterogeneous treatment regions for 100 replications. We compared our algorithm with K-Means, and the results show that K-Means lacks the ability to identify heterogeneous groups in the mediation model(the Appendix A.5 provides detailed procedures and results of the comparative algorithms can be found in Appendix Table A.14,A.15 ,A.16 and Figure A.10). We further compared our method against X-learner, R-learner, and TARNet in the absence of mediators (Appendix A.9), the clustering method comparsion can be found in Appendix A.16.
3.3 Sensitivity and visualization of the M-learner
To assess the robustness of the M-learner, we systematically evaluated its sensitivity to sample size, noise level, number of clusters, projection dimension, projection technique, and clustering algorithm.
@@ 3.3 Sensitivity and visualization of the M-learner
To assess the performance of the M-learner under varying sample sizes, we additionally compare the results to those obtained with a sample size of $500$, while holding all other conditions constant. In Appendix A.4, the Table A.7-A.11 and the Figure A.5-A.9 show the detailed results. The results indicate that, in the absence of mediators, the M-learner is capable of accurately identifying heterogeneous treatment regions in the simple scenario, even with a reduced sample size. However, its performance deteriorates substantially in the complex scenario, suggesting that the identification of treatment effect heterogeneity under more complex conditions requires larger sample sizes. Notably, the M-learner continues to perform well in detecting non-heterogeneous scenarios, even when the available sample size is limited.
@@
In the presence of a mediator, the M-learner remains effective in identifying heterogeneous treatment regions mediated by the mediator, even with a reduced sample size. In the complex scenario, the performance of the RF-based learner declines significantly, whereas the performance degradation of the XGB-based learner is relatively limited. This suggests that, under limited sample sizes, XGB serves as a more robust base learner. The M-learner also demonstrates stable performance; however, accurately estimating complex relationships among $M$, $X$, and $Y$ still requires a sufficiently large sample size to reliably identify heterogeneous treatment effects.
@@
Comparisons of ECDFs across varying sample sizes and scenarios reveal that, in the simple scenario, the ECDFs for sample sizes of $1000$ and $500$ exhibit minimal divergence. In contrast, more pronounced differences are observed in the complex scenario. Notably, an ECDF curve positioned closer to the upper-left corner reflects stronger treatment effect heterogeneity, whereas a curve nearer to the lower-right corner indicates weaker heterogeneity. A comprehensive comparison across different learners, scenarios, and sample sizes is presented in Appendix Figure A.6-A.9.
@@
The visualization results of the M-learner are provided in Appendix A.10 for reference. Further sensitivity analysis experiments can be found in the Appendix A.6.